If Cu(OH)2 has Ksp = 1.6 x 1019, what is the molar solubility of Cu(OH)2?

a) 5.1 × 1010 M

b) 6.4 × 107 M

c) 2.7 × 1011 M

d) 1.7 × 1010 M

e) 3.4 × 107 M
Answer #1


the Ksp of the Cu(OH)2 is as follows

Cu(OH)2 ---------------------> Cu2+ + 2 OH-

I                                              0           0

C                                            x            2x

Ksp = [Cu2+] [OH-]2

Ksp = (x)(2x)2

Ksp = 4x3

=> 1.6*10-19 = 4x3

=> x = 3.4199*10-7 M = molar solubility

answer => e = 3.4*10-7 M